OpenStudy (billyjean):

a diameter of a circle has endpoints p(-10 -2) and q(4 6).B.find the radius .If your answer is not an integer,express it in radical form.

OpenStudy (anonymous):

find the length using these points

OpenStudy (anonymous):

r^2=[(y1-y2)^2+(x1-x2)^2]

OpenStudy (billyjean):

im a little bit confused

OpenStudy (anonymous):

r=sqrt[(6-(-2))^2+(4-(-10))^2] r=sqrt[8^2+14^2]

OpenStudy (billyjean):

thats the answer

OpenStudy (billyjean):

the square root of [8^2+14^2]=2sqrt65

OpenStudy (anonymous):

r-sqrt[64+196] r=sqrt(260)

OpenStudy (billyjean):

thats the answer

OpenStudy (anonymous):

260/4= 65 r=2sqrt(65)

OpenStudy (billyjean):

yes or no is that the answer

OpenStudy (anonymous):

the answer is r=2sqrt(65)